OpenStudy (rayford):

What are the solutions of 2x2 + 8x = –26?

OpenStudy (cwrw238):

= x^2 + 4x + 13 = 0 you can use quadratic formula to solve this

OpenStudy (campbell_st):

rewrite it 2x^2 + 8x + 26 = 0 divide by 2 x^2 + 4x + 13 = 0 you'll need the general quadratic formula or completing the square to solve it

OpenStudy (rayford):

so is it 2+6i?

OpenStudy (campbell_st):

isn't it 2 plus or minus 6i

OpenStudy (rayford):

im sorry yes.. am i correct?

OpenStudy (campbell_st):

yes.... I think so... I only did it quickly as mental arithmetic

OpenStudy (rayford):

i really hope i am right -_-

OpenStudy (rayford):

crap can someone check this for me? cause i realy dont want to get this wrong..

OpenStudy (anonymous):

It would be \[-2 ± 3i\] \[x = \frac{-b ± \sqrt{b^2 - 4ac}}{2a} \implies \frac{-4 ± \sqrt{-36}}{2} \implies -2 ± 3i\]

OpenStudy (rayford):

thank you very much

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

!!! the correct answer is −2 ± 3i !!!